OpenStudy (anonymous):

The cost of renting a compact car from a Rent-A-Car company is 100 per day plus 0.50 per mile driven. In addition, the cost of car insurance is 3.0 per rental. Assuming a linear relationship [a] Write the equation for the cost of renting the car for a day in terms of the number of miles driven. [b] The total cost if the car is rented for a day and it is used for 180 miles.

OpenStudy (anonymous):

based on your question, we have to disregard the car insurance, since it only ask for cost of renting the car for a day in terms of the number of miles driven. am i ryt?

OpenStudy (anonymous):

am not sure but i thinh it should be included

Directrix (directrix):

What does this mean "In addition, the cost of car insurance is 3.0 per rental." Should there be a % symbol after the 3 as in 3% of the rental cost?

OpenStudy (dumbcow):

c= .5m +103 c(180) = 90+103 = 193 not cheap :)

OpenStudy (anonymous):

there is no percentage sign......directrix

OpenStudy (anonymous):

dumbcow......can u pls explain

OpenStudy (anonymous):

representation; r= number of day m=number of mile traveled \[cost=100r+.50m+3\] that's the equation i got.

Directrix (directrix):

3.0 per rental -- $3.00 per day for insurance?

OpenStudy (dumbcow):

well the fixed cost is 103 for day rental and insurance then for each mile you are charged 50 cents

OpenStudy (anonymous):

to find the cost given 180 miles for a day cost= 100(1) + .50(180) + 3 = 100 + 90 + 3 = 193, not bad. . .

OpenStudy (anonymous):

is it clear to you now evelyn?

OpenStudy (anonymous):

yes dear and tanx

OpenStudy (anonymous):

yw (lol, thanks for the dear. just kidding)

OpenStudy (anonymous):

hahahahahahha

OpenStudy (anonymous):

lol... but, anyway thanks for the medal. .

OpenStudy (anonymous):

you welcome